OpenStudy (anonymous):

1 over (b - 1) + 3 over (1 - b) can someone explain how to do this problem please? :)

OpenStudy (anonymous):

Multiply the second fraction by -1/(-1) Then you will have 2 fractions with the same denominator and you can add numerators (1-3)/(b-1)=-2/(b-1) or 2/(1-b)

OpenStudy (anonymous):

1/(b-1) + 3/(1-b) =1/(b-1) - 3/(b-1) = -2/(b-1)

OpenStudy (anonymous):

=2/(1-b)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you are welcome